OpenStudy (rayray1234):

Cindy has 30 meters of fencing. She plans to fence a rectangular dog run that is 4 meters wide. How long can the run be made?

OpenStudy (zpupster):



OpenStudy (zpupster):

perimeter = L + L + w + w P=2L + 2w

OpenStudy (zpupster):

30 meters of fencing so our perimeter can be 30 our equation is the: 30 = 2L + 2(4) 30 = 2L + 8 now we subtract 8 from both sides 30 - 8 = 2L + 8-8 22 = 2L and dived both sides by 2 22/2 = 2L/2 11=L do a check 30 = 11 + 11 + 4 + 4